OpenStudy (j2lie):

How to graph y=x^2-2x-5?

OpenStudy (j2lie):

@whpalmer4

OpenStudy (whpalmer4):

Same way you graph nearly any function. Pick some values of \(x\), calculate the corresponding values of \(y\), plot the points on your graph, draw a smooth curve through them. I'd suggest \(x = -3\) through \(x = 5\) would be an appropriate range

OpenStudy (whpalmer4):

I say that because of the position of the vertex.

OpenStudy (j2lie):

i tried but they don't work. it's very frustrating.

OpenStudy (anonymous):

try using a graphing calculator if your stuck

OpenStudy (whpalmer4):

sure it does. x = -3, y = (-3)^2 - 2(-3) -5 = 9+6-5 = 10 x = -2, y = (-2)^2 - 2(-2) -5 = 4 + 4 - 5 = 3 x = -1, y = (-1)^2 - 2(-1) - 5 = 1 + 2 - 5 = -2 x = 0, y = (0)^2 - 2(0) - 5 = -5 x = 1, y = (1)^2 - 2(1) - 5 = 1-2-5 = -6 x = 2, y = (2)^2 - 2(2) - 5 = 4 - 4 - 5 = -5 x = 3, y = (3)^2 - 2(3) - 5 = 9 - 6 - 5 = -2 x = 4, y = (4)^2 - 2(4) - 5 = 16 - 8 - 5 = 3 x = 5, y = (5)^2 - 2(5) - 5 = 25 - 10 - 5 = 10 plot those points.

OpenStudy (j2lie):

you're a genius. thank you.

OpenStudy (j2lie):

I just only need five.